Circular No. 7069

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1940 AB G. Kulin (1940, BZ 22, 17) remarked that 1940 AB (originally an-

nounced in BZ 22, 11 and on RI 2073) appeared to be a short-period comet, and he computed an orbit with e = 0.45. The object was given the cometary designations 1940a and 1939 VIII, but subsequent attempts to recover the object failed. Acting on advice from the Konkoly Observatory, B. G. Marsden (MPC 6815) computed a main-belt asteroidal orbit for the object (with assumed e = 0.05), calling it again 1940 AB, and he omitted the entry 1939 VIII from the Catalogue of Cometary Orbits, already in the 1972 edition. The object was also excluded from consideration as a comet when the new cometary designation system was introduced in 1994. Very recently, A. Doppler (MPEC 1998-Y10) made the asteroidal identification 1988 RZ_{4} = 1998 KD_{53}, to which G. V. Williams (ibid.) has added further identifications, including 1940 AB, which is therefore now a proven main-belt minor planet having e = 0.09, a = 3.16 AU, i = 14^{o}.

1996 FG_{3}

S. Mottola, Deutsches Zentrum für Luft und Raumfahrt, Berlin, and F. Lahulla, Observatorio Astronómico, Madrid, report: “Our photometric observations of this Apollo object (MPEC 1996-F07, 1997-B05, 1998-Y04, etc.) taken on Dec. 13.0, 14.0 and 15.0 UT with the Max-Planck-Institut 1.23-m telescope at Calar Alto reveal a complex lightcurve with an amplitude of 0.25 mag. Two incommensurate frequencies appear clearly in the Fourier spectrum: f_{1} at 2.97/day and f_{2} at 13.3/day. The same periodicities are also present in our data taken with the 0.60-m Bochum telescope at La Silla over ten nights of observations during the discovery apparition. These features are compatible with an excited, nonprincipal-axis rotation. We encourage observations to characterize the rotational state.”

1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Dec. 17 4^{h}22.^{m}79 +31^{o}30.^{′}4 0.137 1.115 161.^{o}5 16.^{o}2 15.3
19 4 19.90 +30 17.5 0.150 1.126 159.5 17.8 15.6
21 4 17.75 +29 15.4 0.164 1.137 157.4 19.4 15.8
23 4 16.21 +28 22.3 0.177 1.147 155.3 21.0 16.1
25 4 15.18 +27 36.5 0.191 1.157 153.1 22.6 16.3
27 4 14.58 +26 56.9 0.206 1.168 151.0 24.1 16.5
29 4 14.34 +26 22.5 0.220 1.178 149.0 25.5 16.7
31 4 14.43 +25 52.5 0.235 1.187 147.0 26.8 16.9
Jan. 2 4 14.79 +25 26.4 0.250 1.197 145.0 28.1 17.1
1998 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden